### v1

# Automated Defect Detection for Mass-Produced Electronic Components Based on YOLO Object Detection Models

Abstract
Since the defect detection of conventional industry components is time-consuming and labor-intensive, it leads to a significant burden on quality inspection personnel and makes it difficult to manage product quality. In this paper, we propose an automated defect detection system for the dual in-line package (DIP) that is widely used in industry, using digital camera optics and a deep learning (DL)-based model. The two most common defect categories of DIP are examined: (1) surface defects, and (2) pin-leg defects. However, the lack of defective component images leads to a challenge for detection tasks. To solve this problem, the ConSinGAN is used to generate a suitable-sized dataset for training and testing. Four varieties of the YOLO model are investigated (v3, v4, v7, and v9), both in isolation and with the ConSinGAN augmentation. The proposed YOLOv7 with ConSinGAN is superior to the other YOLO versions in accuracy of 95.50%, detection time of 285 ms, and is far superior to threshold-based approaches. In addition, the supervisory control and data acquisition (SCADA) system is developed, and the associated sensor architecture is described. The proposed automated defect detection can be easily established with numerous types of defects or insufficient defect data.
Index Terms:
Deep learning (DL), You Only Look Once (YOLO), automated defect detection, camera sensor, automated optical inspection (AOI), dual inline package (DIP) switch, ConSinGAN.
I Introduction
Global market trend [1] shows that the global market size of the dual in-line package (DIP) [2] switch achieved USD 400.1 million in 2021. Based on a compound annual growth rate of 3.3% during the forecast period, it is predicted to reach USD 554.35 million by 2031. It exhibits a steady annual growth with approximately 10 million units produced daily. With this huge amount of the manufacturing process, it is an essential issue to manage product quality and yield rates. However, conventional quality management depends on human visual product inspections. Due to the time-consuming and labor-intensive nature of the task, it leads to a significant burden on quality inspection personnel. Therefore, the current trend of Industry 4.0 is to streamline manufacturing industries, converting from labor-intensive work toward deep learning (DL)-based automated production.
Over the past decade, DL [3, 4, 5, 6] has been widely applied to various fields, including image recognition, natural language processing, autonomous driving, and object classification problems. Further, DL has also been used for medical radiography image detection and classification [7, 8, 9, 10], A. Jo and W. Lee, [11] proposed a method of deep learning models for material discrimination and quantitation. S. M. J. Jalali et al., [12] proposed a method of deep reinforcement learning-based ensemble algorithm that integrates optimized deep learning models to minimize wind power forecasting errors for two wind power datasets. In [13], a method of deep learning was proposed by E. Jeong et al. for heterogeneous inference parallelization.
Several works [14, 15, 16] use DL to reduce labor resource consumption and promote production line automation. In [14], a dual-weighted analysis algorithm was proposed by K. Qiu et al. for metal defect detection. Y. Chen et al[15] focused on improving the metal punching position using fiber Bragg grating (FBG) sensors and identifying the impact source location through a hyperbolic localization model. In [16], A. Albanese et al. developed a system using a micro-camera control unit (MCU) for automatic DL imaging and detection. Several studies [17, 18, 19, 20] use DL and threshold-based methods to detect defects. L. Xiao et al[17] utilized wavelet transformations to process image features for surface defect detection. Y. Zhang et al[18] formulated defect detection as a classified problem, which could then be resolved via support vector machines (SVM) and random forests (RF). S. Tian et al[19] proposed a complementary adversarial network-driven surface defect detection (CASDD) framework to identify different types of texture defects automatically and accurately. Y. Peng et al[20] proposed a method to detect surface defects on electric distribution cabinets using SVM and Gaussian distribution modeling threshold techniques. A. Dairi et al[21] proposed a novel stereo-vision method that merges a deep encoder with $k$ nearest neighbor (KNN) for anomaly detection to detect obstacles in a road environment. In  [22], A. Dairi et al. proposed a deep encoder with a one-class support vector machine (OCSVM) for obstacle detection systems. Table I summarizes the applications and objects of the methods above.
TABLE I: Comparison of defect detection methods
Methods Model / Algorithm Detected objective Goal
DL CNN Test components [16] Surface defect
Test components [15] Object positioning
Steel [19] Surface defect
IPCNN Car [17] Oil leak detection
KNN Obstacle [21, 22] Obstacle detection
OCSVM
Threshold
Wavelet multiscale
analysis
Tire [18] Surface defect
Gaussian distribution
modeling,
SVM
Electric distribution
cabinet [20]
Surface defect
Dual-weighted
analysis algorithm
Metal parts [14] Surface defect
However, these studies overlooked several key aspects: 1) Real-time detection: Most studies focus on performance in terms of accuracy, and neglect the detection speed performance in the practical application. 2) Computational limitations: It is crucial to lighten the DL model to fit hardware constraints, rather than increasing its size indefinitely for computational capability. 3) Insufficient dataset: The lack of datasets leads to a challenge for detection tasks.
Therefore, we design the lightweight DL model to fit hardware limitations and real-time detection capability. From the review of the image detection models [23, 24, 25, 26], we selected anchor-based one-stage models, specifically the You Only Look Once (YOLO) [27, 28, 29, 30, 31, 32, 33].
YOLO is a neural network that has shown superior performance in terms of both real-time operation and accuracy. YOLO uses the feature pyramid network (FPN) framework enhancement for iterative updates to achieve better results until convergence. YOLO has demonstrated commendable performance in various high-demand real-time scenarios such as public face recognition, mask detection, self-driving car systems, automated manufacturing systems, surface defect detection, etc. Several studies [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46] suggested using YOLO for the dynamic detection of metal surface defects [34, 35, 37, 38]. The authors of [36] applied YOLO to real-time detection for traffic signs. The authors of [43] performed a visual detection of autonomous robots by YOLO. Furthermore, improved YOLO models were used to detect surface defects in solar cells in [40] and [44]. YOLO has also been used for real-time detection of ships [39, 41] and vehicles [42, 45, 46].
To the best of our knowledge, this work is the first study to investigate DL-based automated DIP image detection. A quality-control defect detection method is proposed for each side of this hexagonal component. The proposed system is not a stand-alone system but is integrated with automated manufacturing production line processes. Therefore, the detection time of the automated production output from the preceding workstation must be considered to ensure that detection is completed within a specified time frame.
On the other hand, the primary challenge of automated image processing is the scarcity of defective product samples. To enhance the limited amount of samples, a generative adversarial network (GAN) model [47] called ConSinGAN [48] is used to increase the input image quantity for data argumentation. The ConSinGAN can develop a model based on a single image and effectively simulate the defect characteristics in the generated image. It greatly facilitates the training process of detection models. The other GANs, such as deep convolutional GAN (DCGAN) [49] or Wasserstein GAN (WGAN) [50], require more images for training. The main contributions of this paper are described as follows:
  • The proposed defective DIP detection system is designed based on the YOLO model to improve the quality of the inspection.
  • To enhance the performance of the YOLO model, ConSinGAN is used to generate the amount of DIP image data for the training phase.
  • To validate the effectiveness, we develop a practical production line and SCADA interface to reflect the performance in practical applications.
  • To compare the different YOLO versions, we evaluate the performances in terms of accuracy and detection time.
  • Compared with threshold-based detection, the proposed YOLO models with ConSinGAN reveal a reduction in production detection time by 909-948 ms, where the YOLOv7 with ConSinGAN achieves an accuracy of 95.50%.
  • Unlike previous studies that only focused on detection models, we implement the proposed detection in a practical automated mechanism, while considering the detection time for performance measurement.
The remainder of this paper is organized as follows. Sec. II presents the architecture of the automated defect detection system, the types of DIP defects, and threshold-based image detection for baseline comparison. Sec. III details the data pre-processing, YOLO models, and performance metrics. Sec. IV compares the simulation results among the threshold-based detection, DL-based YOLOv3, v4, v7, and v9 models with/without ConSinGAN for DIP defect detection. Sec. V summarizes the conclusions and the future works.
II Automated Defect Detection System
The system architecture is shown in Fig. 1 and includes three parts: 1) Control system: Comprised of a personal computer (PC) and the programmable logic controller (PLC) devices. The PC is used to interact with the imaging equipment and establish the SCADA, which allows for integration with the DL model and data analysis. The PLCs are used to connect the mechanical equipment, control various actions, issue action completion signals, and provide a secondary confirmation count. 2) Imaging equipment: industrial cameras, centrifugal lenses, and various light source equipment. The imaging equipment interacts and feeds data to the PC via an Ethernet cable (RJ45). 3) Mechanical equipment: The human-machine interface of the PLC includes pneumatic clamps, electromagnetic push rods, and solenoid valves.
Figure 1: System architecture diagram.
II-A Automated Mechanism
The proposed system is designed to inspect all six sides of the workpiece. Each side is equipped with multiple Automated Optical Inspections (AOI) at different locations to minimize interference from various light sources on different surfaces. A 3D schematic diagram of the proposed system architecture is shown in Figure 2. Fig. 2 shows the position and orientation of the AOI and the design of the experimental fixture. It establishes the positioning of lenses and light sources at different angles to inspect the long and short sides of a rectangular workpiece. The detected DIP is placed inversely at the production line. Thus, the DIP top-side detection uses the camera to capture images from bottom to top, and the DIP bottom-side detection captures images from top to bottom. The back- and front- sides of DIP are shot using two sets of cameras from the left and right. Side-to-side imaging is facilitated by rotating the device, allowing cameras on both sides to capture images. The longer sides are the up, down, back, and front parts of the detected DIP. The shorter sides are the left and right of the detected DIP. To examine the longer and shorter surfaces of rectangular DIP, the experiment required two sets of cameras and lenses, each with a specific depth of field. Long-side cameras use a low depth of field, and short-side cameras use a high depth of field. Tables  II and  III list the specific parameters of the camera lens set.
Figure 2: 3D simulation schematic.
TABLE II: Low depth of field camera and lens specifications
BASLER-acA4096-30 µm (Camera) SPO-200I-4M (Len)
Specification Parameter Specification Parameter
Pixel size 3.45 µm Depth of field 4.5
[…]
TABLE III: High depth of field camera and lens specifications
BASLER-acA3800-14 µm (Camera) OPTO-TC23048(Len)
Specification Parameter Specification Parameter
Pixel size 1.67 µm Depth of field 20.0 mm
[…]
The size of each workpiece is approximately 21.7 mm pm\ pm0.5 mm on the long side and 9.4 mm pm\ pm0.4 mm on the short side. The resolving power of the image equipment is calculated by
$RP=\frac{PS}{MF},$ (1)
where $RP$ is the resolving power, $PS$ is the pixel size, and $MF$ is the magnification. The field of view (FOV) is calculated by
${\rm FOV}=RP\times IR,$ (2)
where $IR$ is the image resolution.
II-B Types of Defective DIP
A normal DIP switch is shown in Fig. 3. For defective DIPs, there are two categories: surface and pin-leg. The surface defects include three types: surface overflow, surface scratches, and surface contamination. In addition, misaligned pin-legs are also a type of defect. The above four defective DIPs are shown in Fig. 4.
The sizes of the defect feature are as follows: Fig. 4(a) is approximately 100 to 300 pixels, Fig. 4(b) is approximately 100 to 500 pixels, Fig. 4(c) is approximately 100 to 300 pixels, and Fig. 4(d) is 220 to 250 pixels per pin. To enhance the contrast degree of the feature, the images in Fig. 4(a), 4(b),  4(c), and  4(d) can use the red green blue (RGB) channel subtraction, such as red (R) minus green (G) and G minus blue (B) to enhance the features of the image.
Figure 3: Different sides of Normal DIP, (a) front, (b) back, (c) bottom, (d) top, (e) right, and (f) left sides.
Figure 4: Four types of defective DIP, (a) surface glue overflow, (b) surface scratches, (c) surface dirt, and (d) bent pins.
II-C Threshold-Based Image Detection
For a baseline comparison, we introduce a threshold-based image detection. First, it performs binary classification of the original image using a preset threshold to obtain the featured image. The feature extraction is determined by
$b(x,y)=\begin{cases}255,\mathrm{MinGray}\leq g(x,y)\leq\mathrm{MaxGray},\\ 0,\ \rm otherwise,\end{cases}$ (3)
where $b(x,y)$ represents the pixel of the featured image based on the binary decisions, $g(x,y)$ is the pixel of the original image, $\rm MinGray$ is the minimum threshold, and $\rm MaxGray$ is the maximum threshold.
Since the light intensity corresponding to each camera varies with the six sides of the DIP surface, the threshold values $\rm MinGray$ and $\rm MaxGray$ are adjustable. Whether each side of the DIP surface is defeated or not is determined by
$DR(z)=\begin{cases}\mathrm{defective\;DIP},\ TB\leq o(z)\\ \mathrm{Normal\;DIP},\ TB>o(z)>0,\end{cases}$ (4)
where $DR(z)$ represents the threshold-based detected result, $TB$ represents the preset threshold, $z$ denotes the featured image, and $o(\cdot)$ is the output image of feature extraction. The features of defective DIP are shown in Fig. 5.
Figure 5: The defective features from threshold-based detection.
III DL-based YOLO Model and ConSinGAN
Fig. 6 shows the workflow of the proposed detection. To select the outstanding YOLO version, we compare different YOLO versions in terms of accuracy, floating-point operations (FLOPs), and the number of parameters. To enhance the accuracy of the YOLO models, we augment the data set by using ConSinGAN [48] during the data pre-processing. The ConSinGAN modifies the original images using rotation, translation, flipping, scaling, Gaussian noise, etc, for data augmentation.
Figure 6: The flow chart of the proposed defective DIP detection.
III-A Data Pre-Processing
III-A1 Data Augmentation
To augment the defective DIP image dataset, various techniques are used: flipping, mirroring, brightness adjustment, median filtering, noise, and Gaussian blur, as depicted in Fig. 7. Data augmentation is crucial to improve the trained model and enhance the discrimination.
Figure 7: Image augmentation. (a) brightness adjustment, (b) mirroring, (c) median filtering, (d) flipping, (e) noise, and (f) Gaussian blur.
III-A2 ConSinGAN
The predecessor of ConSinGAN, SinGAN [51], introduces a GAN that is trained on a single image for tasks such as unconditional image generation and harmonization. The Concept of the ConSinGAN model is shown in Fig. 8.
Figure 8: The concept of the ConSinGAN.
ConSinGAN is trained using a multi-stage and multi-resolution approach with the lowest resolution (e.g., 25 × 25 pixels) in the first stage. The next stage increases the layers of the neural network and image resolution. At each stage, the layers of the neural network in all the previous stages are frozen, and only the current added layers of the neural network are trained. Different from the traditional GAN, it uses a multi-generator, and the number of stages increases the number of generators. The traditional GAN propagates feature maps from one generator stage to the next, which leads to a negative effect on learning performance. The training of ConSinGAN is limited to one stage at a specific time to prohibit interactions between different stages. If all stages are end-to-end training, it causes overfitting in the single-image case, and the network collapses. It means that the relationship between the receptive field and the size of the generated image decreases as the number of stages increases. At higher resolutions, the discriminator focuses on the texture of the image. At lower resolutions, the discriminator concerns the global layout of the image. At a given stage $i$, ConSinGAN optimizes the sum of an adversarial and a reconstruction loss, and initializes the discriminator with the weights of the previous stage i1i-1 at all stages as follows:
$\min_{G_{i}}\max_{D_{i}}\mathcal{L}_{adv}(G_{i},D_{i})=\alpha\mathcal{L}_{rec}(G_{i}),$ (5)
where $\mathcal{L}_{adv}(G_{i},D_{i})$ is the adversarial loss that is the same with WGAN [50], the reconstruction loss $L_{rec}(G_{i})$ is used to improve the training stability ($\alpha$ is set to 10 in our experiments) as follows:
[…]
At the given resolution of stage $i$, we input $(s_{0})$ into the generator $G_{i}(s_{0})$. The input $(s_{0})$ is a downsampling version of the original image $(s_{i})$.
III-B YOLO Learning Series
YOLO is adopted as a DL-based object detection algorithm. Compared with traditional object detection algorithms, the YOLO model can quickly perform object detection and classification within a single neural network. The YOLO architecture is a convolutional neural network (CNN) that includes multiple convolutional and fully connected layers, segments an image into S × S grids, and then detects multiple objects in each grid. In addition, the selection criteria of the different YOLO versions for DIP testing were based on their floating-point operations (FLOPs) and mAP. To compare different YOLO versions, we list the performance in terms of FLOPs, and of YOLO v3, v4, v5, v6, v7, v8, and v9, as shown in Table IV. YOLOv7 [29] is chosen as the proposed DIP testing model. In our simulation, YOLOv3 [27], YOLOv4  [28], and the current newest YOLOv9 [33] are considered for comparison. As shown in Table VI, the YOLOv7 slightly outperforms YOLOv9 in terms of accuracy but is slightly slower. YOLOv5 [30] and YOLOv6 [31] have similar FLOP performance to YOLOv7, but their mAP is less than that of YOLOv7. YOLOv8 has higher computation complexity than YOLOv7 in terms of FLOP by 63%; however, the mAP improved by only 0.1%. Therefore, YOLOv5, v6, and v8 are excluded from our main simulations.
TABLE IV: Comparison of the YOLO series
Model #Param (M) FLOPs mAP0.5
YOLOv3 [27] 62 1457B 57.9%
YOLOv4 [28] 53 109B 65.7%
YOLOv5-L [30] 46.5 109.1G 67.3%
YOLOv6-M [31] 34.9 85.8G 66.4%
YOLOv7 [29] 36.9 104.7G 69.7%
YOLOv8-L [32] 43.7 165.2G 69.8%
YOLOv9-C [33] 25.3 102.1G 70.2%
To further compare the YOLOv3, v4, v7, and v9 models, we introduce the upgraded parts of these models in the following:
III-B1 YOLOv3 [27]
The most significant improvement of YOLOv3 is adopting the multiscale feature extraction, also known as a feature pyramid network (FPN). In addition, YOLOv3 uses binary cross-entropy as the loss function, instead of the mean square error of YOLOv2 [53]. YOLOv3 uses darknet-53 to replace darknet-19 of YOLOv2 and increase depth and accuracy, but it sacrifices the detection speed.
III-B2 YOLOv4 [28]
First, YOLOv4 uses CSPDarknet53 to establish the primary backbone neural network, making it a lightweight model while maintaining high accuracy. Second, YOLOv4 adopts spatial pyramid pooling [54] and path aggregation network [55] in the neck section to combine feature maps at different scales and detect different sizes of objects. Finally, the YOLOv4 employs CIoU-loss to evaluate the distance between the targeted and the predicted boxes and improve the accuracy.
III-B3 YOLOv7 [29]
YOLOv7 [29] and YOLOv4 [28] were developed by H.-Y. M. Liao, et al. The backbone of YOLOv7 is redesigned based on v4. The backbone is used to extract the features from the input images. Unlike YOLOv5 and v6, YOLOv7 is not pre-trained on ImageNet. YOLOv7 has several new architectures with improved performances in terms of accuracy and detection speed. The architecture of YOLOv7 is shown in Fig. 9
Figure 9: The YOLOv7 network architecture and block description.
In Fig. 9, YOLOv7 uses extended ELAN (E-ELAN), which develops the original ELAN architecture through expanding, shuffling, and merging. These approaches allow the network to learn more diverse features without disrupting the original gradient paths. The head part of YOLOv7, which is used for object detection and localization, continues to utilize the head network structure of YOLOR [56]. The CBS block includes the convolutional layer, batch normalization, and the Silu activation function. The CBM block consists of the convolutional layer, batch normalization, and the Sigmoid activation function. REP block represents the RepVGG. The difference between MP-1 and MP-2 is the output size within the block. The difference between E-ELAN and ELAN-W is the output size within the block. The main improvements of YOLOv7 are as follows:
  • In RepConv, the identity connection (id) disrupts the residual in ResNet and the concatenation in DenseNet. However, RepConv needs to retain features that provide greater gradient diversity for different feature maps. A novel weight parameterization approach is proposed in [57] for RepConv to replace id. In Fig. 10, the planned reparametrized convolution designed for ResNet is called RepConvN [29]. Since PlainNet in Fig. 10(a) inherently lacks residual or concatenation characteristics, RepConv in Fig. 10(b) can be applied directly. Since ResNet in Fig. 10(c) inherently has residual characteristics, RepConv in Fig. 10(d) leads to a decrease in accuracy. Based on the above reasons, YOLOv7 uses RepConvN (RepConv without id) to design a network architecture.
Figure 10: The proposed RepConvN using (a) PlainNet, (b) RepPlainNet, (c) ResNet, and (d) RepResNet.
  • Supervised learning is commonly used to train deep networks, where the weights of shallower layers are trained using an auxiliary mechanism to improve the accuracy of YOLOv7. In addition, YOLOv7 introduces a novel dynamic label assignment strategy that employs a coarse-to-fine guide to enhance feature learning. The final output head is called the lead head, and the head used for auxiliary training is called the auxiliary head. Fig. 11 depicts an object detector equipped with an auxiliary head.
Figure 11: YOLO model with auxiliary head.
  • By optimizing ELAN, as shown in Fig. 12(a), the shortest and longest gradient paths are controlled to allow deeper layers to learn and converge more effectively. However, stacking more CNN layers in ELAN may disrupt stable learning. Therefore, YOLOv7 proposes extended ELAN (E-ELAN) in Fig. 12(b) to maintain gradient paths while continuously enhancing the learning performance using expanding, shuffling, and merging of cardinalities.
Figure 12: ELAN optimization schematic for (a) ELAN, and (b) E-ELAN.
III-B4 YOLOv9 [33]
YOLOv9 introduces two novel techniques, including programmable gradient information (PGI) and generalized efficient layer aggregation network (GELAN). The descriptions of PGI and GELAN are as follows:
  • The PGI theory involves redesigning a technique called the auxiliary reversible branch that allows the YOLOv9 to generate reliable gradient information during the training phase and pass it to the main branch.
  • GELAN combines the functions of CSPNet and ELAN networks. It splits the input of each computational unit into two parts: one part is directly passed to the output, while the other part is transformed by the computational unit. Then, these two parts of computation are combined to form the final output. Thus, GELAN integrates information between computational units while enhancing flexibility and generalization.
The proposed model consists of two parts: data augmentation with ConSinGAN and training the YOLO model. The limitations of the YOLO model are as follows:
  • Computational capability: Depending on the constraints of hardware, it is necessary to use a lightweight YOLO model, which may sacrifice accuracy.
  • Training cost: To achieve accurate detection, the YOLO model requires a large amount of labeled data for training because collecting and labeling sufficient training data is a time-consuming and costly task.
  • External factors: In the image, additional obstacles or illumination changes may reduce the accuracy of the YOLO model.
  • Size of defect features: If the defect features are tiny, the single-stage detection is inferior to multi-stage detection.
III-C Performance Evaluation
The leave-one-out cross-validation (LOOCV) is used to measure the performance of the proposed DIP detection system. The testing set includes one subject’s data, while the training set uses the remaining subjects’ data. LOOCV is repeated $k$ times, where $k$ is the total number of subjects, until all subjects have been used as the testing set. By using $k$-folds, the average performance can be obtained.
In our experiments, we adopt a confusion matrix as performance measurement, and five evaluation metrics, precision ($PRE$), recall ($REC$), F1-Score ($F1$), false positive rate ($FPR$), and true negative rate ($TNR$) as shown from Eq. (7) to (11) to measure the detection performance.
$PRE=\rm\frac{TP}{TP+FP},$ (7)
$REC=\rm\frac{TP}{TP+FN},$ (8)
$F1=2\times\frac{PRE\cdot REC}{PRE+REC},$ (9)
$FPR=\rm\frac{FP}{FP+TN},$ (10)
$TNR=\rm\frac{TN}{FP+TN},$ (11)
where TP (true positive), TN (true negative), FP (false positive), and FN (false negative) represent the cases: the defect and normal DIPs labeled as such are correctly recognized, the normal DIPs labeled as defective DIPs are misclassified, and the defective DIPs labeled as normal DIPs are misclassified, respectively.
The performance evaluation of detection depends on the intersection over the union (IoU), and mean average precision (mAP), where mAP0.5 means the threshold of YOLO is set to 0.5. All model training is conducted on a PC with an Intel 13th Gen I7-13700K CPU (3.40 GHz), NVidia RTX4080 graphics card with 16 GB of dedicated memory, running on a 64-bit Windows 10 operating system. The training and testing phase for the models is executed in an Anaconda (a Python derivative) environment, and the SCADA interface is integrated using Visual Studio 2019.
IV Experimental Results
In our experiments, the nine kinds of models are specified as follows: 1) Threshold-based detection, 2) YOLOv3, 3) YOLOv4, 4) YOLOv7, 5) YOLOv9, 6) YOLOv3 with ConSinGAN, 7) YOLOv4 with ConSinGAN, 8) YOLOv7 with ConSinGAN, 9) YOLOv9 with ConSinGAN. The hyperparameter settings of these models are as follows:
  • The threshold-based detection: Based on Eq. (4), $TB$ is assigned a constant value of 100.
  • The proposed YOLOv3, v4, v7, and v9 models with/without ConSinGAN: Batches = 32, images = $\times416$, learning rate = 0.001, and maximum batches = 10000.
The hyperparameter settings for the generative adversarial network ConSinGAN are as follows: Learning rate = 0.1, the number of trained stages = 10, and the original images are captured using the industrial camera described in Sec. II. The total number of original images is 672. The ConSinGAN selects images with significant defective features for image augmentation. The total number of generated images is 3,183.
IV-A DIP images dataset
IV-A1 ConSinGAN Augmentation
Initially, GAN and DCGAN were used to augment data. However, the training performances of GAN and DCGAN were poor because of the limited number of defective image samples. Therefore, ConSinGAN is used, which generates similar images from a smaller number of images for data augmentation. The original images are shown in Fig. 13, and the augmented results are shown in Fig. 14. The amount of each side of the defective DIP in the generated dataset is shown in Table V.
TABLE V: Statistics of ConSinGAN-generated images.
Type of defect Front Back Top bottom Left Right Total
Surface defect 340 228 209 183 310 346 1616
Pin defect 321 216 201 177 325 327 1567
Figure 13: Original images of defective DIP, (a) top, (b) back, (c) right, (d) front, (e) bottom, and (f) left sides.
Figure 14: Images generated by ConSinGAN, (a) top, (b) back, (c) right, (d) front, (e) bottom, and (f) left sides.
IV-B Performances of YOLO with/without ConSinGAN
The YOLO model is trained for each aspect of the generated image data. These 3,183 images are divided into two categories: surface defects (surface overflow, surface scratches, and surface dirt) and pin defects (pin misalignment). We compare YOLOv3, v4, v7, and v9 with/without the ConSinGAN augmented sets to identify the optimal model. The training set is conducted on 80% of the dataset, i.e., 2,547 images. The validation set is adopted on the other 10% (318 images). The remaining 10% (318 images) are used for the testing set. Within this dataset, we define two distinct defect types that include: 1) Surface defects account for approximately 50.7% of defects, and 2) Misaligned pins account for 49.3%. Table VI demonstrates the performances of YOLOv3, v4, v7, and v9 with/without the ConSinGAN in terms of average precision (mAP) rate, $PRE$, RECREC, $F1$, FPRFPR, and $TNR$ evaluations of all six YOLO models.
TABLE VI: Performance Comparison of YOLO models.
Model mAP0.5 PREPRE RECREC $F1$ FPRFPR TNRTNR Detection time (ms)
YOLOv3 66.2% 73.2% 51.5% 61.8% 35.0% 64.9% 322
YOLOv4 65.4% 80.1% 61.4% 68.9% 35.2% 64.8% 314
YOLOv7 65.1% 84.2% 58.6% 68.1% 34.3% 65.6% 297
YOLOv9 74.5% 79.2% 66.4% 72.8% 40.3% 59.7% 279
YOLOv3 with ConSinGAN 91.9% 92.8% 91.4% 92.4% 8.9% 91.1% 321
YOLOv4 with ConSinGAN 89.1% 91.3% 92.6% 91.5% 10.7% 89.2% 310
YOLOv7 with ConSinGAN 95.5% 94.9% 90.7% 93.4% 3.7% 96.4% 285
YOLOv9 with ConSinGAN 90.5% 87.7% 82.9% 85.0% 9.8% 90.2% 278
As shown in Table VI, 1) With ConSinGAN, YOLOv3, v4, v7, and v9 outperform the models without ConSinGAN in terms of accuracy metrics due to data augmentation. By data augmentation, the YOLOv3, v4, v7, and v9 have been enhancing the performance in the training phase because of the increase in the amount of training data. 2) YOLOv7 with ConSinGAN achieves outstanding performance on accuracy metrics and outperforms the current newest YOLOv9 with ConSinGAN. The proposed YOLOv7 with ConSinGAN model achieves an accuracy in terms of mAP0.5 = 95.5% and outperforms the other models. However, the detection speed of YOLOv9 with ConSinGAN outperforms the other models and is slightly better than YOLOv7 with ConSinGAN. The detection time of each model is approximately 300 ms. Therefore, we select the YOLOv7 model in a DIP image detection system on the SCADA interface for the user, as illustrated in Fig. 15. In addition, ConSinGAN is only used to augment the dataset. Thus, it does not impact the complexity of YOLO models.
Figure 15: The SCADA interface.
IV-C Performances of Threshold-based Detection
For baseline comparison, we introduce the performance evaluation of threshold-based detection. Table VII presents the number of images, accuracy, and detection time from different sides of DIP by threshold-based detection. Each workpiece is inspected from all six sides. The number of images is the same as in the YOLO experiments. The threshold value is set to 0.5. The accuracy is calculated by
$A=\frac{DT-F}{DT}\times 100\%,$ (12)
where $A$ represents the accuracy, $DT$ is the total number of detections, $F$ is the number of detection errors.
Table VII shows the average accuracy is $87.81\%$ on six-sided DIP image detection. The total detection time is 3.807 seconds, which matches the detection time. Our experiment involves an actual production line machine, divided into six-sided detectors that are moved synchronously. Therefore, the detection time is determined by the station with the longest detection time. The proposed YOLOv7 with ConSinGAN model outperforms the threshold-based system, achieving an accuracy of 95.5%.
IV-D Comparison of Detection Speed
Fig. 16 demonstrates the detection time of the YOLO models and shows a significantly higher detection speed than the threshold-based system. The threshold-based method needs to reset the parameters for every six-sided DIP image. However, YOLO models only need one setting, because the training and testing stages are both performed using six-sided DIP images as input. Different from threshold-based image processing, YOLO models do not require parameter adjustments to different sides of DIP to obtain accurate results. Thus, there are performance gaps between YOLO models and the threshold-based method. From Fig. 16, it is feasible to replace the threshold-based image-processing defect detection rate with the proposed YOLOv7 with the ConSinGAN approach.
TABLE VII: Performance of the threshold-based detection
Detected side Test number Accuracy Detection time (ms)
Left 635 86.30% 290
Top 410 86.58% 391
Right 673 88.85% 359
bottom 360 92.22% 815
Back 444 90.54% 721
Front 661 84.56% 1231
Total 3183 87.81% 3807
Figure 16: Comparison of detection speed.
V Conclusion
In this paper, our goal is to establish an AOI system for defective DIP detection to improve the quality of production and save human efforts. We use the SCADA interface to integrate image detection and automated mechanism hardware for a large number of DIP inspections. For the image detection model, we use threshold-based detection and DL-based YOLO models. The threshold-based detection outperforms the YOLO models in terms of accuracy. However, the threshold-based detection is time-consuming. Thus, we adopt ConSinGAN to augment the dataset for YOLO training and improve the accuracy. With ConSinGAN, YOLOv7 achieves an accuracy of $95.5\%$, outperforming the accuracy of $87.81\%$ by threshold-based detection. The detection times of different YOLO models are 285 to 322 ms, which outperforms 1231 ms of the threshold-based detection. In summary, ConSinGAN effectively enhances the accuracy of YOLO models while maintaining the outstanding detection time. For future works, we summarize as follows:
  • Applying future novel models to improve the detection model network.
  • Increasing the variety of workpieces for the automated detection system, such as rotating DIP and surface mount devices, which are mass-produced electronic components.
  • Expanding the SCADA system to cover additional production lines, aiming for smart factories and AIoT integration.